OpenStudy (anonymous):

F=9/5C+32 solve for C Please Help me show the work, I know the answer is, C=F-32/9

jimthompson5910 (jim_thompson5910):

First you subtract 32 from both sides. Then you multiply both sides by 5/9 to isolate C \[\large F = \frac{9}{5}C + 32\] \[\large F-32 = \frac{9}{5}C + 32-32\] \[\large F-32 = \frac{9}{5}C\] \[\large \frac{9}{5}C = F-32\] \[\large \frac{5}{9}*\frac{9}{5}C = \frac{5}{9}(F-32)\] \[\large C = \frac{5}{9}(F-32)\]

OpenStudy (anonymous):

thank you!!!!

jimthompson5910 (jim_thompson5910):

you're welcome